A metal has Er=9, wo=7×10^-15 rad s^-1, v=2×10^12 s. Calculate the skin depth for microwave of 4 cm vacuum wavelength
1
Expert's answer
2019-04-26T11:17:29-0400

The skin depth is


"\delta=\sqrt{\frac{\rho}{\pi f\mu}}=\sqrt{\frac{\rho\lambda}{\pi c\mu}}=\sqrt{\frac{1.68\cdot 10^{-8}\cdot 0.04}{3.14\cdot 3\cdot 10^8\cdot 1.26\cdot 10^{-6}}}=0.75\space\mu \text{m}."